It is known that 10% of bolts manufactured by a factory are defective. If a random sample of 100 bolts is chosen at random from a day's production, construct the sampling distribution of (1) the number of defective bolts, (2) the proportion of defective bolts.
1
Expert's answer
2020-05-12T17:31:02-0400

1) N=n*p=100*0.1=10

2) sampling distribution of a sample proportion

"\mu_\rho=p=0.1,"

"\sigma_\rho=\sqrt{\frac{p(1-p)}{n}}" =0.03.